# BiTrajDiff: Bidirectional Trajectory Generation with Diffusion Models for Offline Reinforcement Learning

### v3

@@ 5.4 Effectiveness under $n$-step TD Estimator
We employ the $n$-step TD estimator to update the critic networks in IQL and TD3BC under the original dataset and the DA-augmented data, and further compare performance across different values of $n$ as shown in Figure 5. The $n$-step TD estimator (de2018multi) enables the discovery of high-return solutions by looking ahead from more distant states, making performance improvement a natural indicator of the quality of DA-generated datasets. As shown in Figure 5, all three DA methods consistently outperform the base method, suggesting that the generated trajectories effectively uncover more rewarding states. However, as $n$ increases, the performance inevitably declines due to the growing variance in value estimation. Meanwhile, for DA methods, this effect is compounded by the accumulated mismatching errors introduced by the inverse dynamics model, which negatively impact TD estimation. In particular, our method consistently achieves the highest performance across all tested settings and maintains its advantage over RTDiff and Diffstitch as the $n$-step horizon increases. These results verify the effectiveness of BiTrajDiff for offline RL algorithms under $n$-step TD estimator. By stitching forward-future and backward-history trajectories, BiTrajDiff obtain a superior capability to synthesize high-quality trajectories compared to existing DA methods.
@@ 5.5 Visualization
To evaluate the accuracy and diversity robustness of the BiTrajDiff framework, we compare its generated trajectories with those produced by single-direction forward and backward diffusion models. Following lu2023synthetic, we employ two quantitative metrics to assess a generated trajectory $\hat{\tau}$: (1) Dynamic Error: $\mathcal{E}_{\text{Dyn}}(\hat{\tau})=\sum_{t}\|\hat{s}_{t+1}-s_{t+1}\|_{2}$, which measures trajectory accuracy by summing the mean squared error between the predicted next states $\hat{s}_{t+1}$ and the ground truth $s_{t+1}$; (2) L2 Distance: $\mathcal{E}_{\text{L2D}}(\hat{\tau})=\min_{\tau\in\mathcal{D}}\sum_{t}\|\hat{s}_{t}-s_{t}\|_{2}$, defined as the minimal L2 distance between the generated trajectory and any trajectory of equal length in $\mathcal{D}$, capturing trajectory diversity. As shown in Figure 6, while forward and backward diffusion models achieve low $\mathcal{E}_{\text{Dyn}}$, they exhibit limited $\mathcal{E}_{\text{L2D}}$, indicating that the generated trajectories closely resemble those in the original dataset and thus offer limited performance gains. In contrast, trajectories generated by BiTrajDiff demonstrate significantly broader marginal distributions in $\mathcal{E}_{\text{L2D}}$ while maintaining competitive $\mathcal{E}_{\text{Dyn}}$, reflecting enhanced behavioral diversity without sacrificing dynamic consistency. Consequently, BiTrajDiff facilitates greater performance gains with the augmented data. These results highlight the ability of BiTrajDiff to generate behaviorally diverse trajectories while preserving accuracy, effectively balancing exploration and reliability.
@@ B.2.1 BiTrajDiff Implementation
In this section, we provide the implementation details of our experiments. We conducted our experiments on a cluster of 4 A100 GPUs. The source code will be made publicly available upon the publication of this paper. Our BiTrajDiff is implemented based on CleanDiffuser (dong2024cleandiffuser), a recent modularized library for diffusion models in decision making. We represent the noise model $\epsilon_{\theta}$ with the 1D DiT backbone (Peebles2022DiT), consisting $2$ transformer blocks with adaptive layer normalization. Each block contains a multi-head self-attention layer followed by a feed-forward MLP. The diffusion timestep embeddings are first encoded via Fourier features and then projected through a two-layer MLP network with 128 hidden units. Similarly, the condition inputs are encoded by a two-layer MLP network with 128 hidden units. For the sampling process, we follow the CleanDiffuser reimplementation of Decision Diffuser (ajay2022conditional), employing the VP-SDE (song2020score) formulation with 20 diffusion steps for sampling. Meanwhile, the inverse dynamics and rewards model are both represented by a two-layer MLP network with 512 hidden units, respectively. During BiTrajDiff training, we employ an Adam optimizer with a learning rate of 2e42e-4 for all networks, using a batch size of $64$ and performing $1e6$ gradient steps in total. The default trajectory horizon $H$ for both training and generation is set to $5$. The ratio $\sigma$ of the augmented dataset is $30\%$. During our BiTrajDiff data augmentation pipeline, the batch size is set to $512$, while the filter number $C_{\text{ood}}$ is $256$ and $C_{\text{greedy}}$ is $64$. As for the target return-to-go $R^{f}$ and $R^{b}$, we set them to be equal in our BiTrajDiff conditional generation. Each task is assigned its own target return, and in our experiments we directly adopt the corresponding hyperparameters from the CleanDiffuser reimplementation of Decision Diffuser.
@@ B.2.2 Offline RL Algorithm Implementation
In our experiments, we evaluate BiTrajDiff against four offline RL algorithms: IQL, TD3BC, CQL, and DT. To enable efficient and large-scale evaluation, we conduct all offline RL experiments using the JAX-based library JAX-CORL (nishimori2024jaxcorl). For fairness, we report re-run results on JAX-CORL with initial datasets as baseline performance, rather than the scores reported in the original papers, to account for discrepancies across codebases. Meanwhile, for each offline RL algorithm, we ensure that the hyperparameters are configured consistently with those reported in the respective original papers.
@@ Appendix C Experimental results on CQL and DT
Source Task  CQL (kumar2020conservative)  DT (chen2021decision)
@@
Task  CQL (kumar2020conservative)  DT (chen2021decision)
@@ Appendix D Experimental results with the direction of diffusion generation
Source Task  IQL (kostrikov2021offline)  TD3BC (fujimoto2021minimalist)
@@
Source Task  CQL (kumar2020conservative)  DT (chen2021decision)
@@ Appendix E Experimental results with trajectory filters
Source Task  IQL (kostrikov2021offline)  TD3BC (fujimoto2021minimalist)
@@ Appendix G Experimental results with $n$-step TD Estimator
Length of n-step TD Estimator($n$)   IQL (kostrikov2021offline)  TD3BC (fujimoto2021minimalist)
@@
Length of n-step TD Estimator($n$)   IQL (kostrikov2021offline)  TD3BC (fujimoto2021minimalist)

### v4

@@ 5.4 Effectiveness under $n$-step TD Estimator
We employ the $n$-step TD estimator to update the critic networks in IQL and TD3BC under the original dataset and the DA-augmented data, and further compare performance across different values of $n$ as shown in Figure 5. The $n$-step TD estimator (De Asis et al., 2018) enables the discovery of high-return solutions by looking ahead from more distant states, making performance improvement a natural indicator of the quality of DA-generated datasets. As shown in Figure 5, all three DA methods consistently outperform the base method, suggesting that the generated trajectories effectively uncover more rewarding states. However, as $n$ increases, the performance inevitably declines due to the growing variance in value estimation. Meanwhile, for DA methods, this effect is compounded by the accumulated mismatching errors introduced by the inverse dynamics model, which negatively impact TD estimation. In particular, our method consistently achieves the highest performance across all tested settings and maintains its advantage over RTDiff and Diffstitch as the $n$-step horizon increases. These results verify the effectiveness of BiTrajDiff for offline RL algorithms under $n$-step TD estimator. By stitching forward-future and backward-history trajectories, BiTrajDiff obtain a superior capability to synthesize high-quality trajectories compared to existing DA methods.
@@ 5.5 Visualization
To evaluate the accuracy and diversity robustness of the BiTrajDiff framework, we compare its generated trajectories with those produced by single-direction forward and backward diffusion models. Following Lu et al. (2023), we employ two quantitative metrics to assess a generated trajectory $\hat{\tau}$: (1) Dynamic Error: $\mathcal{E}_{\text{Dyn}}(\hat{\tau})=\sum_{t}\|\hat{s}_{t+1}-s_{t+1}\|_{2}$, which measures trajectory accuracy by summing the mean squared error between the predicted next states $\hat{s}_{t+1}$ and the ground truth $s_{t+1}$; (2) L2 Distance: $\mathcal{E}_{\text{L2D}}(\hat{\tau})=\min_{\tau\in\mathcal{D}}\sum_{t}\|\hat{s}_{t}-s_{t}\|_{2}$, defined as the minimal L2 distance between the generated trajectory and any trajectory of equal length in $\mathcal{D}$, capturing trajectory diversity. As shown in Figure 6, while forward and backward diffusion models achieve low $\mathcal{E}_{\text{Dyn}}$, they exhibit limited $\mathcal{E}_{\text{L2D}}$, indicating that the generated trajectories closely resemble those in the original dataset and thus offer limited performance gains. In contrast, trajectories generated by BiTrajDiff demonstrate significantly broader marginal distributions in $\mathcal{E}_{\text{L2D}}$ while maintaining competitive $\mathcal{E}_{\text{Dyn}}$, reflecting enhanced behavioral diversity without sacrificing dynamic consistency. Consequently, BiTrajDiff facilitates greater performance gains with the augmented data. These results highlight the ability of BiTrajDiff to generate behaviorally diverse trajectories while preserving accuracy, effectively balancing exploration and reliability.
@@ B.2.1 BiTrajDiff Implementation
In this section, we provide the implementation details of our experiments. We conducted our experiments on a cluster of 4 A100 GPUs. The source code will be made publicly available upon the publication of this paper. Our BiTrajDiff is implemented based on CleanDiffuser (Dong et al., 2024), a recent modularized library for diffusion models in decision making. We represent the noise model $\epsilon_{\theta}$ with the 1D DiT backbone (Peebles and Xie, 2022), consisting $2$ transformer blocks with adaptive layer normalization. Each block contains a multi-head self-attention layer followed by a feed-forward MLP. The diffusion timestep embeddings are first encoded via Fourier features and then projected through a two-layer MLP network with 128 hidden units. Similarly, the condition inputs are encoded by a two-layer MLP network with 128 hidden units. For the sampling process, we follow the CleanDiffuser reimplementation of Decision Diffuser (Ajay et al., 2022), employing the VP-SDE (Song et al., 2020) formulation with 20 diffusion steps for sampling. Meanwhile, the inverse dynamics and rewards model are both represented by a two-layer MLP network with 512 hidden units, respectively. During BiTrajDiff training, we employ an Adam optimizer with a learning rate of 2e42e-4 for all networks, using a batch size of $64$ and performing $1e6$ gradient steps in total. The default trajectory horizon $H$ for both training and generation is set to $5$. The ratio $\sigma$ of the augmented dataset is $30\%$. During our BiTrajDiff data augmentation pipeline, the batch size is set to $512$, while the filter number $C_{\text{ood}}$ is $256$ and $C_{\text{greedy}}$ is $64$. As for the target return-to-go $R^{f}$ and $R^{b}$, we set them to be equal in our BiTrajDiff conditional generation. Each task is assigned its own target return, and in our experiments we directly adopt the corresponding hyperparameters from the CleanDiffuser reimplementation of Decision Diffuser.
@@ B.2.2 Offline RL Algorithm Implementation
In our experiments, we evaluate BiTrajDiff against four offline RL algorithms: IQL, TD3BC, CQL, and DT. To enable efficient and large-scale evaluation, we conduct all offline RL experiments using the JAX-based library JAX-CORL (Nishimori, 2024). For fairness, we report re-run results on JAX-CORL with initial datasets as baseline performance, rather than the scores reported in the original papers, to account for discrepancies across codebases. Meanwhile, for each offline RL algorithm, we ensure that the hyperparameters are configured consistently with those reported in the respective original papers.
@@ Appendix C Experimental results on CQL and DT
Source Task  CQL (Kumar et al., 2020)  DT (Chen et al., 2021)
@@
Task  CQL (Kumar et al., 2020)  DT (Chen et al., 2021)
@@ Appendix D Experimental results with the direction of diffusion generation
Source Task  IQL (Kostrikov et al., 2021)  TD3BC (Fujimoto and Gu, 2021)
@@
Source Task  CQL (Kumar et al., 2020)  DT (Chen et al., 2021)
@@ Appendix E Experimental results with trajectory filters
Source Task  IQL (Kostrikov et al., 2021)  TD3BC (Fujimoto and Gu, 2021)
@@ Appendix G Experimental results with $n$-step TD Estimator
Length of n-step TD Estimator($n$)   IQL (Kostrikov et al., 2021)  TD3BC (Fujimoto and Gu, 2021)
@@
Length of n-step TD Estimator($n$)   IQL (Kostrikov et al., 2021)  TD3BC (Fujimoto and Gu, 2021)